A length of aluminum wire is connected to a precision 10.00-V power supply, and a current of 0.4583A is precisely measured at 25.0?C. The wire is placed in a new environment of unknown temperature where the measured current is 0.3673A.

What is the unknown temperature?.
Answer #1

Resistance at 25 C

Ro=V/I =10/0.4583

Ro=21.82 ohms

Resistance at Unknown temperatur e T is

R=10/0.3673

R=27.23 ohms

R=Ro[1+alpha*(T-To)]

27.23=21.82[1+0.00429*(T-25)]

T=82.75 C